### v1

# Group class operations and homological conditions

Abstract.
Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ and Talelli’s operation $\Phi$ can be often used to formally enlarge the class of available examples of groups that satisfy certain homological conditions. In this paper, we employ this enlargement technique regarding two specific homological conditions. We thereby demonstrate the abundance of groups that (a) have virtually Gorenstein group algebras, as defined by Beligiannis and Reiten, and (b) satisfy Moore’s conjecture on the relation between projectivity and relative projectivity, that was studied by Aljadeff and Meir.
0. Introduction
The study of group actions on spaces is very useful in cohomological group theory. It offers a geometric intuition that often clarifies purely algebraic results and also leads to important constructions that help us understand fine aspects of the group structure. In order to examine certain homological finiteness conditions on groups and explore their relation to complete cohomology, Kropholler has introduced in [30] the ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups, building upon earlier work by Ikenaga [28]. These groups are constructed hierarchically, starting with the class of finite groups, and constitute a very big class; it does require some effort to find groups which are not contained therein. It turns out that the hierarchical definition of these groups provides an effective tool for extending homological properties of finite groups to a much wider class. From the extensive literature on this topic, we only mention the early prototypical works [30], [7] and [17]. More generally, we may apply Kropholler’s construction for any group class $\mathfrak{C}$ and obtain a group class ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ (which is, in principle, much bigger that $\mathfrak{C}$). Another example of a group class operation is provided by Talelli’s construction of groups of type $\Phi$; these groups were defined in [36], in order to construct groups that admit a finite dimensional model for the classifying space for proper actions. Even though the motivation originates from geometric considerations, the definition of groups of type $\Phi$ is purely algebraic: It is based on the need to describe modules of finite projective dimension over the group, in terms of the finiteness of the projective dimension of their restriction to finite subgroups. More generally, starting with any class of groups $\mathfrak{C}$, Talelli’s construction provides us with a group class $\Phi\mathfrak{C}$; see also [32]. We may replace the projective dimension with either the flat or the injective dimension and obtain the analogous group class operations $\Phi_{flat}$ and $\Phi_{inj}$. An iteration of Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ and Talelli’s operation $\Phi$ (in some of its three forms) often enlarges the class of available examples of groups that satisfy appropriate homological conditions.
Examples of this kind of homological conditions can be found in the realm of Gorenstein homological algebra, the relative homological theory which is based upon the classes of Gorenstein projective, Gorenstein flat and Gorenstein injective modules introduced in [23] and [25]. This theory has been developed rapidly and found interesting applications in algebraic geometry, representation theory and cohomological group theory. Even though important advances have been made recently by Šaroch and Št’ovíček [35], there are still many properties of the classical homological algebra notions that are not universally available for their Gorenstein homological algebra counterparts. The technique of using the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ was employed in [22], in order to exhibit a very big class of groups that is convenient to work with, as far as these properties are concerned, in the sense that these properties are satisfied by the Gorenstein modules over the corresponding group algebras. In the present work, we provide two more instances of such an enlargement of the class of examples of groups that satisfy homological restrictions. The first of these examples deals with the notion of virtually Gorenstein algebras. This notion was introduced for Artin algebras by Beligiannis and Reiten in [5], in order to study the left-right symmetry of Gorenstein regularity; cf. [4, Theorem 8.7]. It provides a common generalization of Gorenstein regular algebras and algebras of finite representation type. The implications of the virtually Gorenstein condition to the homotopy (or the stable) categories of Gorenstein projective and Gorenstein injective modules are analysed in [3]; see also [11] and [37]. This notion was extended to arbitrary rings in [18]; it is proved therein that all weakly Gorenstein regular rings are virtually Gorenstein. Having fixed the commutative ring $k$, we may consider the class $\mathfrak{W}$ that consists of those groups $G$ for which $kG$ is weakly Gorenstein regular. If the class $\mathfrak{W}$ is non-empty, i.e. if $k$ is itself weakly Gorenstein regular, then $\mathfrak{W}$ contains all finite groups. For any $\mathfrak{W}$-group $G$, [18, Theorem A] implies that $kG$ is virtually Gorenstein. We now enlarge $\mathfrak{W}$ and let $\overline{\mathfrak{W}}$ be the smallest group class which contains $\mathfrak{W}$ and is both ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Then, $\overline{\mathfrak{W}}$-groups admit a hierarchical description à la Kropholler, starting with $\mathfrak{W}$ and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi_{inj}\mathfrak{C}$; this process is analysed in §$\S$1.IV. The enlargement of the group class of available examples that is advocated in this paper, is illustrated for the virtually Gorenstein property by the following result.
Theorem A. If $G$ is a $\overline{\mathfrak{W}}$-group, then $kG$ is virtually Gorenstein.
This result is proved in §$\S 4$, using the description of the $\mbox{Ext}^{1}$-orthogonal classes of Gorenstein projective or Gorenstein injective modules over group algebras of groups that are obtained by performing Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ (see also [29]) or Talelli’s operation $\Phi$ to certain group classes. The relation between these operations and $\mbox{Ext}^{1}$-orthogonality is analysed in §$\S 3$.
Another example of a homological condition that is well-behaved with respect to the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ is provided by the relation between projectivity and relative projectivity for modules over group algebras. As an attempt to generalize the crucial step in the proof of Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [10, Chapter VIII, §$\S$3], Moore has conjectured that a module over the group algebra of a group $G$ must be necessarily projective, provided that its restriction to a subgroup $H\subseteq G$ of finite index (that satisfies an additional necessary group-theoretic condition) is projective. Chouinard’s work [14] implies that all finite groups satisfy Moore’s conjecture. More generally, Aljadeff and Meir [1] have proved that Moore’s conjecture is satisfied by all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups. Building upon their work, in the present paper we improve the latter result as follows. Let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed. As before, $\overline{\mathfrak{F}}$-groups admit a hierarchical description starting with the class $\mathfrak{F}$ of finite groups and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi\mathfrak{C}\cup\Phi_{flat}\mathfrak{C}$. The following result is proved in §$\S 5$.
Theorem B. If $G$ is an $\overline{\mathfrak{F}}$-group, then $G$ satisfies Moore’s conjecture over $k$.
Here is a brief description of the contents of the paper: Following the preliminary Section 1, in Section 2, we fix the convenient universe of groups within which all Gorenstein homological algebra notions behave smoothly. In Section 3, we build on Kendall’s work [29] and describe the orthogonal modules with respect to the $\mbox{Ext}^{1}$-pairing of the Gorenstein classes over the group algebras of groups that are obtained by performing Kropholler’s or Talelli’s operation to group classes within the universe of groups described in Section 2. As an application of this description, we prove Theorem A in Section 4. Finally, in Section 5 we discuss Moore’s conjecture and combine the results by Aljadeff and Meir [1] with certain simple observations regarding the relation between projectivity and relative projectivity to prove Theorem B.
Notations and terminology. Let $k$ be a commutative ring and $G$ a group. We shall denote by ${\tt Proj}(kG)$, ${\tt Flat}(kG)$ and ${\tt Inj}(kG)$ the classes of projective, flat and injective $kG$-modules respectively. We shall also denote by $\mathcal{P}(kG)$, $\mathcal{F}(kG)$ and $\mathcal{I}(kG)$ the classes of $kG$-modules that appear as kernels of acyclic complexes of projective, flat or injective $kG$-modules respectively. Finally, if $M$ is a $kG$-module, then its character (Pontryagin dual) module is the $kG$-module $DM=\mbox{Hom}_{\mathbb{Z}}(M,\mathbb{Q}/\mathbb{Z})$ of all additive maps from $M$ to $\mathbb{Q}/\mathbb{Z}$.
1. Preliminaries
In this preliminary section, we record certain basic notions that are used in the sequel. These notions concern restriction and (co-)induction of modules over group algebras, basic properties of Gorenstein modules, cotorsion pairs in module categories and the hierarchical construction of the closure of group classes under certain operations.
I. Restriction and (co-)induction. Let $k$ be a commutative ring, $G$ a group and $H\subseteq G$ a subgroup. Restricting the action of $G$ on a $kG$-module $M$ to the subgroup $H$, we obtain the $kH$-module $\mbox{res}_{H}^{G}M$. In this way, we define the restriction functor $\mbox{res}_{H}^{G}$ from the category of $kG$-modules to that of $kH$-modules. This functor has both a left and a right adjoint. Its left adjoint is the induction functor $\mbox{ind}_{H}^{G}$; for any $kH$-module $N$ the induced $kG$-module $\mbox{ind}_{H}^{G}N$ is equal to $kG\otimes_{kH}N$, with $G$ acting on the left of $kG$. The right adjoint of restriction is the coinduction functor $\mbox{coind}_{H}^{G}$. Here, for any $kH$-module $N$ the coinduced $kG$-module $\mbox{coind}_{H}^{G}N$ is equal to $\mbox{Hom}_{kH}(kG,N)$, with $G$ acting on the right of $kG$. Any $kH$-module $N$ is contained as a direct summand in both $kH$-modules $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}N$ and $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}N$. For more details, the reader is referred to [10, Chapter III, §$\S$3]. Deriving the two adjunction isomorphisms, we obtain the Echmann-Shapiro isomorphisms: If $M$ is a $kG$-module and $N$ a $kH$-module, there are natural identifications
$\mbox{Ext}_{kH}^{n}\!\left(N,\mbox{res}_{H}^{G}M\right)\!\simeq\mbox{Ext}_{kG}^{n}\!\left(\mbox{ind}_{H}^{G}N,M\right)\;\;\mbox{and}\;\;\;\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N\right)\!\simeq\mbox{Ext}_{kG}^{n}\!\left(M,\mbox{coind}_{H}^{G}N\right)$
for all $n\geq 0$. We also note that, in the special case where $H$ has finite index in $G$, the functors $\mbox{ind}_{H}^{G}$ and $\mbox{coind}_{H}^{G}$ are naturally isomorphic; cf. [10, Chapter III, Proposition 5.9].
II. Gorenstein modules. An acyclic complex of projective modules over a ring $R$ is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(\_\!\_,P)$ for any projective module $P$. A module $M$ is Gorenstein projective if it is a cokernel of a totally acyclic complex of projective modules; we denote by ${\tt GProj}(R)$ the class of these modules. Dually, an acyclic complex of injective modules is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(I,\_\!\_)$ for any injective module $I$. A module $N$ is Gorenstein injective if it is a kernel of a totally acyclic complex of injective modules; we denote by ${\tt GInj}(R)$ the class of these modules. An acyclic complex of flat modules is called totally acyclic if it remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$. A module $L$ is Gorenstein flat if it is a cokernel of a totally acyclic complex of flat modules; we denote by ${\tt GFlat}(R)$ the class of these modules. The Gorenstein modules defined above were introduced by Enochs, Jenda and Torecillas in [23] and [25]; see also Holm’s paper [27]. A distinguished class of Gorenstein flat modules has been introduced by Šaroch and Št’ovíček in [35]: A module $K$ is projectively coresolved Gorenstein flat if it is a cokernel of an acyclic complex of projective modules, which remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$; the class of these modules is denoted by ${\tt PGF}(R)$. As shown in [35, Theorem 4.4], projectively coresolved Gorenstein flat modules are Gorenstein projective, i.e. ${\tt PGF}(R)\subseteq{\tt GProj}(R)$.
[…]
III. Cotorsion pairs. If $R$ is a ring, then the $\mbox{Ext}^{1}$-pairing induces an orthogonality relation between modules. In this way, if S is a class of modules, the left orthogonal ${}^{\perp}{\tt S}$ of ${\tt S}$ is the class consisting of those modules $M$, for which $\mbox{Ext}^{1}_{R}(M,S)=0$ for all $S\in{\tt S}$. Analogously, the right orthogonal ${\tt S}^{\perp}$ of S is the class consisting of those modules $N$, for which $\mbox{Ext}^{1}_{R}(S,N)=0$ for all $S\in{\tt S}$. If ${\tt U},{\tt V}$ are two module classes, we say that $({\tt U},{\tt V})$ is a cotorsion pair (cf. [24]) if ${\tt U}=\!\,^{\perp}{\tt V}$ and ${\tt U}^{\perp}={\tt V}$. The cotorsion pair is hereditary if $\mbox{Ext}^{i}_{R}(U,V)=0$ for all $i>0$ and all $U\in{\tt U}$ and $V\in{\tt V}$; it is complete if for any module $M$ there exist short exact sequences
$0\longrightarrow V\longrightarrow U\longrightarrow M\longrightarrow 0\;\;\;\mbox{and }\;\;\;0\longrightarrow M\longrightarrow V^{\prime}\longrightarrow U^{\prime}\longrightarrow 0,$
where $U,U^{\prime}\in{\tt U}$ and $V,V^{\prime}\in{\tt V}$. We say that the cotorsion pair $({\tt U},{\tt V})$ is cogenerated by a class of modules ${\tt S}\subseteq{\tt U}$ if ${\tt S}^{\perp}={\tt U}^{\perp}$, so that ${\tt V}={\tt S}^{\perp}$ and ${\tt U}={{}^{\perp}}\!\left({\tt S}^{\perp}\right)$. As shown in [19], any cotorsion pair which is cogenerated by a set of modules is necessarily complete.
Non-trivial examples of cotorsion pairs are provided by Gorenstein modules. As shown in [35], there are complete hereditary cotorsion pairs $\left({\tt PGF}(R),{\tt PGF}(R)^{\perp}\right)$, $\left({\tt GFlat}(R),{\tt GFlat}(R)^{\perp}\right)$ and $\left({}^{\perp}{\tt GInj}(R),{\tt GInj}(R)\right)$ for any ring $R$. The situation regarding Gorenstein projective modules is more subtle: It is shown in [15] that $\left({\tt GProj}(R),{\tt GProj}(R)^{\perp}\right)$ is a hereditary cotorsion pair as well, but its completeness appears to depend upon set-theoretic assumptions.
IV. Continuous operations and the closure of group classes. An operation $\mathcal{O}$ on group classes assigns a group class $\mathcal{O}\mathfrak{C}$ to each group class $\mathfrak{C}$, in such a way that the following two properties are satisfied:
(i) $\mathfrak{C}\subseteq\mathcal{O}\mathfrak{C}$ for any group class $\mathfrak{C}$ and
(ii) $\mathcal{O}\mathfrak{C}_{1}\subseteq\mathcal{O}\mathfrak{C}_{2}$ for any two group classes $\mathfrak{C}_{1},\mathfrak{C}_{2}$ with $\mathfrak{C}_{1}\subseteq\mathfrak{C}_{2}$.
We say that the operation $\mathcal{O}$ is continuous if it also has the following property:
(iii) Assume that $\mathfrak{C}_{\alpha}$ is a group class defined for each ordinal number $\alpha$, such that $\mathfrak{C}_{\alpha}\subseteq\mathfrak{C}_{\beta}$ for all $\alpha\leq\beta$, and let $\mathfrak{C}$ be the class consisting of those groups which are contained in $\mathfrak{C}_{\alpha}$ for some ordinal number $\alpha$. Then, the group class $\mathcal{O}\mathfrak{C}$ is the class consisting of those groups which are contained in $\mathcal{O}\mathfrak{C}_{\alpha}$ for some $\alpha$.
A group class $\mathfrak{C}$ is said to be closed under an operation $\mathcal{O}$ if $\mathcal{O}\mathfrak{C}=\mathfrak{C}$. The $\mathcal{O}$-closure of a group class $\mathfrak{C}$ is the smallest $\mathcal{O}$-closed group class containing $\mathfrak{C}$; it is the class consisting of those groups $G$ which are contained in $\mathfrak{D}$ for any $\mathcal{O}$-closed class $\mathfrak{D}$ that contains $\mathfrak{C}$. If the operation $\mathcal{O}$ is continuous, then there is an alternative hierarchical (bottom-to-top) description of the $\mathcal{O}$-closure of a group class $\mathfrak{C}$, that we shall now explicit. For any group class $\mathfrak{C}$, we define the classes $\mathfrak{C}_{\alpha}$ for any ordinal number $\alpha$, using transfinite induction, as follows:
(a) $\mathfrak{C}_{0}=\mathfrak{C}$,
(b) $\mathfrak{C}_{\alpha+1}=\mathcal{O}\mathfrak{C}_{\alpha}$ for any ordinal $\alpha$ and
(c) $\mathfrak{C}_{\alpha}=\bigcup_{\beta<\alpha}\mathfrak{C}_{\beta}$ for any limit ordinal $\alpha$.
Let $\overline{\mathfrak{C}}$ be the class consisting of those groups which are contained in $\mathfrak{C}_{\alpha}$, for some ordinal $\alpha$. It is easily seen that $\overline{\mathfrak{C}}$ is the smallest $\mathcal{O}$-closed group class that contains $\mathfrak{C}$; in other words, $\overline{\mathfrak{C}}$ is the $\mathcal{O}$-closure of $\mathfrak{C}$.
If $\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{n}$ are (continuous) operations on group classes, then we may consider the (continuous) operation $\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{2}\cup\ldots\cup\mathcal{O}_{n}$, which is defined by letting $\mathcal{O}\mathfrak{C}=\mathcal{O}_{1}\mathfrak{C}\cup\mathcal{O}_{2}\mathfrak{C}\cup\ldots\cup\mathcal{O}_{n}\mathfrak{C}$ for any group class $\mathfrak{C}$. The $\mathcal{O}$-closure of $\mathfrak{C}$ is the smallest group class which contains $\mathfrak{C}$ and is closed under all of the operations $\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{n}$.
We are interested in the continuous operation ${\scriptstyle{{\bf LH}}}$ on group classes that was introduced by Kropholler in [30]. If $\mathfrak{C}$ is any group class, then we define for each ordinal $\alpha$ the group class ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{C}$, using transfinite induction, as follows: First of all, we let ${\scriptstyle{{\bf H}}}_{0}\mathfrak{C}=\mathfrak{C}$. For an ordinal $\alpha>0$, the class ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{C}$ consists of those groups $G$ which admit a cellular action on a finite dimensional contractible CW-complex $X$, in such a way that each isotropy subgroup of the action belongs to ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{C}$ for some ordinal $\beta<\alpha$ (that may depend upon the particular cell). If $G,X$ are as above, $k$ is a commutative ring and $M$ is a $kG$-module, then the cellular chain complex of $X$ induces two exact sequences of $kG$-modules
$0\longrightarrow M_{r}\longrightarrow\cdots\longrightarrow M_{1}\longrightarrow M_{0}\longrightarrow M\longrightarrow 0$
and
$0\longrightarrow M\longrightarrow M^{0}\longrightarrow M^{1}\longrightarrow\cdots\longrightarrow M^{r}\longrightarrow 0,$
where $r$ is the dimension of $X$ and the $M_{i}$’s (resp. the $M^{i}$’s) are direct sums (resp. direct products) of $kG$-modules of the form $\mbox{ind}_{H}^{G}\mbox{res}_{H}^{G}M$ (resp. $\mbox{coind}_{H}^{G}\mbox{res}_{H}^{G}M$), for some ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{C}$-subgroup $H$ of $G$ with $\beta<\alpha$. We say that a group belongs to ${\scriptstyle{{\bf H}}}\mathfrak{C}$ if it belongs to ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{C}$ for some $\alpha$. Then, the class ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups $G$, all of whose finitely generated subgroups $H$ are contained in some ${\scriptstyle{{\bf H}}}\mathfrak{C}$-subgroup $K\subseteq G$. The continuity of the operation ${\scriptstyle{{\bf LH}}}$ is proved in [22, Lemma 1.3]. If the class $\mathfrak{C}$ is subgroup-closed, then the classes ${\scriptstyle{{\bf H}}}\mathfrak{C}$ and ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ are subgroup-closed as well; in this case, ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups all of whose finitely generated subgroups are ${\scriptstyle{{\bf H}}}\mathfrak{C}$-groups.
The concept of groups of type $\Phi$, that were introduced by Talelli in [36], corresponds to another continuous operation, applied to the class $\mathfrak{F}$ of finite groups. To be more precise, if $k$ is a commutative ring and $\mathfrak{C}$ is a class of groups, then we define $\Phi\mathfrak{C}$ as the class consisting of those groups $G$, for which the following two conditions are equivalent for any $kG$-module $M$:
(i) $\mbox{pd}_{kG}M<\infty$,
(ii) $\mbox{pd}_{kH}\mbox{res}_{H}^{G}M<\infty$ for any $\mathfrak{C}$-subgroup $H\subseteq G$.
See also [9, Definition 5.6]. The continuity of $\Phi$ is proved in [22, Lemma 1.4]. We shall also consider the version of the operation $\Phi$ for the flat (resp. injective) dimension of modules and define for any group class $\mathfrak{C}$ the class $\Phi_{flat}\mathfrak{C}$ (resp. $\Phi_{inj}\mathfrak{C}$) accordingly.
2. A convenient class of groups for studying Gorenstein modules
There are a few conjectural properties of the class of Gorenstein (projective, flat or injective) modules over group algebras that are only known to hold in particular cases. An example of such a property is the stability under restriction to subgroups: If $M$ is a Gorenstein (projective, flat or injective) $kG$-module and $H\subseteq G$ is a subgroup, then is the restricted $kH$-module $\mbox{res}_{H}^{G}M$ also Gorenstein (projective, flat or injective respectively)? Analogous questions involve the completeness of the Gorenstein projective cotorsion pair, the relation between Gorenstein projective and Gorenstein flat modules and the duality between Gorenstein flat and Gorenstein injective modules.
We fix a commutative ring $k$ and let $\mathfrak{X}=\mathfrak{X}(k)$ be the class consisting of those groups $G$, for which ${\tt PGF}(kG)={\tt GProj}(kG)={\mathcal{P}}(kG)$. In other words, we ask that all cokernels of acyclic complexes of projective $kG$-modules be projectively coresolved Gorenstein flat. We also let $\mathfrak{X}^{\prime}=\mathfrak{X}^{\prime}(k)$ be the class consisting of those groups $G$, for which ${\tt GFlat}(kG)={\mathcal{F}}(kG)$ and $\mathfrak{Y}=\mathfrak{Y}(k)$ be the class consisting of those groups $G$, for which ${\tt GInj}(kG)={\mathcal{I}}(kG)$. In other words, $G$ is an $\mathfrak{X}^{\prime}$-group (resp. a $\mathfrak{Y}$-group) if all cokernels (resp. kernels) of acyclic complexes of flat (resp. injective) $kG$-modules are necessarily Gorenstein flat (resp. Gorenstein injective). These classes have been studied in [22], where the following result is proved.
Theorem 2.1.
The classes $\mathfrak{X}$, $\mathfrak{X}^{\prime}$ and $\mathfrak{Y}$ are subgroup-closed and have the following properties:
(i) $\mathfrak{X}=\mathfrak{X}^{\prime}$; cf. [22, Corollary 2.3].
(ii) $\mathfrak{X}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$, $\Phi$ and $\Phi_{flat}$; cf. [22, Theorem 3.3]
(iii) $\mathfrak{Y}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{inj}$; cf. [22, Theorem 4.1].
For later use, we record the following properties of Gorenstein projective modules over a group ring $kG$. The first three of these properties are valid over all groups, whereas the latter three are valid under additional assumptions:
(1) If $H\subseteq G$ is a subgroup, then induction maps ${\tt GProj}(kH)$ into ${\tt GProj}(kG)$.
(2) For any $n\geq 1$ any Gorenstein projective $kG$-module $M$ is an $n$-th syzygy in a projective resolution of another Gorenstein projective $kG$-module $M^{\prime}$.
(3) The orthogonal class ${\tt GProj}(kG)^{\perp}$ is thick [^1] We recall that a class of modules is thick if it is closed under direct summands, extensions, cokernels of monomorphisms and kernels of epimorphisms. and contains all $kG$-modules of finite projective dimension.
(4) If $H$ is an $\mathfrak{X}$-subgroup of $G$, then restriction maps ${\tt GProj}(kG)$ into ${\tt GProj}(kH)$.
(5) If $G$ is an $\mathfrak{X}$-group, then $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is a complete hereditary cotorsion pair.
(6) If $G$ is an $\mathfrak{X}$-group, then ${\tt GProj}(kG)^{\perp}$ contains all $kG$-modules of finite flat dimension.
Proof. Property (1) is precisely [21, Lemma 2.6(i)], whereas (2) and (3) are immediate consequences of the definition of Gorenstein projectivity; see also [27, Theorem 2.20]. Property (4) follows since restriction maps $\mathcal{P}(kG)$ into $\mathcal{P}(kH)={\tt GProj}(kH)$, whereas (5) and (6) follow since ${\tt GProj}(kG)={\tt PGF}(kG)$, in view of [35, Theorems 4.4 and 4.9]. $\Box$
For completeness, we also list the analogous properties of Gorenstein injective modules:
(1’) If $H\subseteq G$ is a subgroup, then coinduction maps ${\tt GInj}(kH)$ into ${\tt GInj}(kG)$.
(2’) For any $n\geq 1$ any Gorenstein injective $kG$-module $M$ is an $n$-th cosyzygy in an injective resolution of another Gorenstein injective $kG$-module $M^{\prime}$.
(3’) The orthogonal class ${}^{\perp}{\tt GInj}(kG)$ is thick and contains all $kG$-modules of finite injective dimension.
(4’) If $H$ is a $\mathfrak{Y}$-subgroup of $G$, then restriction maps ${\tt GInj}(kG)$ into ${\tt GInj}(kH)$.
(5’) The pair $\left({}^{\perp}{\tt GInj}(kG),{\tt GInj}(kG)\right)$ is a complete hereditary cotorsion pair.
Proof. Property (1’) follows by dualizing the argument in [21, Lemma 2.6(i)], whereas (2) is an immediate consequence of the definition of Gorenstein injective modules. Properties (3’) and (5’) follow from [35, Theorem 5.6] and [27, Theorem 2.22]. Finally, property (4’) follows since restriction maps $\mathcal{I}(kG)$ into $\mathcal{I}(kH)={\tt GInj}(kH)$. $\Box$
The following simple (and widely known) observation is the starting point for the description of the classes which are orthogonal to Gorenstein modules in the next section.
Corollary 2.2.
If $H\subseteq G$ is any subgroup, then restriction maps:
(i) ${\tt GProj}(kG)^{\perp}$ into ${\tt GProj}(kH)^{\perp}$ and
(ii) ${}^{\perp}{\tt GInj}(kG)$ into ${}^{\perp}{\tt GInj}(kH)$.
Proof. Assertion (i) follows from the induction-restriction (Eckmann-Shapiro) isomorphism, in view of property (1) above. Analogously, assertion (ii) follows from the restriction-coinduction (Eckmann-Shapiro) isomorphism, in view of property (1’) above. $\Box$
3. $\mbox{Ext}^{1}$-orthogonality and group class operations
We consider the class $\mathfrak{X}$ defined in the previous section; in view of Theorem 2.1(ii), $\mathfrak{X}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$, $\Phi$ and $\Phi_{flat}$. We note that assertions (i) and (ii) below are essentially due to Kendall [29].
Theorem 3.1.
Let $\mathfrak{X}_{0}$ be a subclass of $\mathfrak{X}$. We consider a group $G$ and a $kG$-module $M$, such that $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
(i) We have $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any ${\scriptstyle{{\bf H}}}\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
(ii) If $\mathfrak{X}_{0}$ is subgroup-closed, then $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
(iii) We have $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\Phi\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
(iv) We have $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\Phi_{flat}\mathfrak{X}$-subgroup $H\subseteq G$.
Proof. (i) We shall proceed by induction on the ordinal $\alpha$, for which $H\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{X}_{0}$. The base case where $\alpha=0$ is precisely our hypothesis on $M$. For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{X}_{0}$-subgroups of $G$ and all $\beta<\alpha$. We consider an ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{X}_{0}$-subgroup $H\subseteq G$ and let $N\in{\tt GProj}(kH)$. Then, as noted in §$\S$1.IV, there exists an exact sequence of $kH$-modules
$0\longrightarrow N_{d}\longrightarrow\ldots\longrightarrow N_{1}\longrightarrow N_{0}\longrightarrow N\longrightarrow 0,$
where $d$ is the dimension of the $H$-CW-complex witnessing that $H\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{X}_{0}$ and each $N_{i}$ is a direct sum of $kH$-modules of the form $\mbox{ind}_{F}^{H}\mbox{res}_{F}^{H}N$, with $F$ an ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{X}_{0}$-subgroup of $H$ for some $\beta<\alpha$. Since ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{X}_{0}\subseteq{\scriptstyle{{\bf H}}}_{\beta}\mathfrak{X}\subseteq{\scriptstyle{{\bf H}}}\mathfrak{X}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{X}=\mathfrak{X}$, we may invoke property (4) in Section 2 and conclude that restriction maps ${\tt GProj}(kH)$ into ${\tt GProj}(kF)$; in particular, $\mbox{res}_{F}^{H}N\in{\tt GProj}(kF)$. As our induction hypothesis implies that $\mbox{res}_{F}^{G}M\in{\tt GProj}(kF)^{\perp}$, we have
$\mbox{Ext}^{n}_{kH}\!\left(\mbox{ind}_{F}^{H}\mbox{res}_{F}^{H}N,\mbox{res}_{H}^{G}M\right)\!=\mbox{Ext}^{n}_{kF}\!\left(\mbox{res}_{F}^{H}N,\mbox{res}_{F}^{G}M\right)\!=0$
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(N_{i},\mbox{res}_{H}^{G}M\right)\!=0$ for all $i$ and all $n>0$. Then, using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>d$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the $d$-th syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
$\mbox{Ext}_{kH}^{1}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=\mbox{Ext}_{kH}^{d+1}\!\left(N^{\prime},\mbox{res}_{H}^{G}M\right)\!=0.$
This completes the inductive step of the proof, taking care of any ${\scriptstyle{{\bf H}}}\mathfrak{X}_{0}$-subgroup $H$ of $G$.
(ii) We now assume that $\mathfrak{X}_{0}$ is subgroup-closed and consider an ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}$-subgroup $H$ of $G$. We proceed by induction on the cardinality $\kappa$ of $H$. If $\kappa\leq\aleph_{0}$, then the ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}$-group $H$ is actually contained in ${\scriptstyle{{\bf H}}}\mathfrak{X}_{0}$ and we are done by (i) above. If $\kappa$ is uncountable, then we may express $H$ as a continuous ascending union of subgroups $(H_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Since the class ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}$ is subgroup-closed, $H_{\lambda}$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}$-group and our induction hypothesis implies that $\mbox{res}^{G}_{H_{\lambda}}M\in{\tt GProj}(kH_{\lambda})^{\perp}$ for all $\lambda$. We fix $N\in{\tt GProj}(kH)$ and note that $\mbox{res}^{H}_{H_{\lambda}}N\in{\tt GProj}(kH_{\lambda})$ for all $\lambda$; this follows since $H_{\lambda}\in{\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{X}=\mathfrak{X}$, in view of property (4) in Section 2. Letting $N_{\lambda}=\mbox{ind}_{H_{\lambda}}^{H}\mbox{res}^{H}_{H_{\lambda}}N$, we compute
$\mbox{Ext}^{n}_{kH}\!\left(N_{\lambda},\mbox{res}_{H}^{G}M\right)\!=\mbox{Ext}^{n}_{kH_{\lambda}}\!\left(\mbox{res}^{H}_{H_{\lambda}}N,\mbox{res}_{H_{\lambda}}^{G}M\right)\!=0$
for all $\lambda$ and $n>0$. The continuous ascending family of subgroups $(H_{\lambda})_{\lambda<\kappa}$ induces a continuous direct system of $kH$-modules $(N_{\lambda})_{\lambda<\kappa}$ with surjective structure maps, whose colimit is $N$. The short exact sequence of $kH$-modules
$0\longrightarrow K_{\lambda}\longrightarrow N_{0}\longrightarrow N_{\lambda}\longrightarrow 0,$
where $K_{\lambda}$ is the kernel of the structure map $N_{0}\longrightarrow N_{\lambda}$, is the $\lambda$-th term of a continuous direct system of short exact sequences. The colimit of the latter direct system of short exact sequences is the short exact sequence of $kH$-modules
(1) $0\longrightarrow K\longrightarrow N_{0}\longrightarrow N\longrightarrow 0$
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $N_{\lambda}\longrightarrow N_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kH}(K_{\lambda+1}/K_{\lambda},\mbox{res}_{H}^{G}M)=0$ for all $\lambda$ and $n>0$. Therefore, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}\!\left(K,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>0$. Then, the short exact sequence (1) above implies that $\mbox{Ext}^{n}_{kH}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GProj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${\tt GProj}(kH)^{\perp}$. Indeed, the Gorenstein projective $kH$-module $N$ is the first syzygy in a projective resolution of another Gorenstein projective $kH$-module $N^{\prime}$ (this is property (2) in Section 2) and hence
$\mbox{Ext}_{kH}^{1}\!\left(N,\mbox{res}_{H}^{G}M\right)\!=\mbox{Ext}_{kH}^{2}\!\left(N^{\prime},\mbox{res}_{H}^{G}M\right)\!=0.$
This completes the proof of (ii).
(iii) We fix a $\Phi\mathfrak{X}_{0}$-subgroup $H\subseteq G$. Since $\Phi\mathfrak{X}_{0}\subseteq\Phi\mathfrak{X}=\mathfrak{X}$, we may invoke the completeness of the cotorsion pair $\left({\tt GProj}(kH),{\tt GProj}(kH)^{\perp}\right)$ (property (5) in Section 2) to obtain a short exact sequence of $kH$-modules
$0\longrightarrow\mbox{res}_{H}^{G}M\longrightarrow L\longrightarrow N\longrightarrow 0,$
where $L\in{\tt GProj}(kH)^{\perp}$ and $N\in{\tt GProj}(kH)$. Then, for any $\mathfrak{X}_{0}$-subgroup $F\subseteq H$ we have an exact sequence of restricted $kF$-modules
$0\longrightarrow\mbox{res}_{F}^{G}M\longrightarrow\mbox{res}_{F}^{H}L\longrightarrow\mbox{res}_{F}^{H}N\longrightarrow 0$
Corollary 2.2(i) implies that $\mbox{res}_{F}^{H}L\in{\tt GProj}(kF)^{\perp}$, whereas property (4) of Section 2 shows that $\mbox{res}_{F}^{H}N\in{\tt GProj}(kF)$. Since $\mbox{res}_{F}^{G}M\in{\tt GProj}(kF)^{\perp}$, in view of our assumption on $M$, the thickness of the class ${\tt GProj}(kF)^{\perp}$ implies that $\mbox{res}_{F}^{H}N$ is also contained in ${\tt GProj}(kF)^{\perp}$. Since ${\tt GProj}(kF)\cap{\tt GProj}(kF)^{\perp}={\tt Proj}(kF)$, it follows that $\mbox{res}_{F}^{H}N$ is a projective $kF$-module. This is the case for any $\mathfrak{X}_{0}$-subgroup $F$ of the $\Phi\mathfrak{X}_{0}$-group $H$ and hence $\mbox{pd}_{kH}N<\infty$. It follows that $N\in{\tt GProj}(kH)^{\perp}$ and hence the thickness of ${\tt GProj}(kH)^{\perp}$ implies that $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$, as needed.
(iv) We fix a $\Phi_{flat}\mathfrak{X}_{0}$-subgroup $H\subseteq G$ and follow verbatim the argument in (iii) above. In the very last sentence of that proof, we use the fact that the class ${\tt GProj}(kH)^{\perp}$ contains all modules of finite flat dimension; the latter claim follows from property (6) in Section 2, since $H\in\Phi_{flat}\mathfrak{X}_{0}\subseteq\Phi_{flat}\mathfrak{X}=\mathfrak{X}$. $\Box$
Corollary 3.2.
Let $\mathfrak{X}_{0}$ be a subgroup-closed subclass of $\mathfrak{X}$ and consider a group $G$, which is contained in ${\scriptstyle{{\bf LH}}}\mathfrak{X}_{0}\cup\Phi\mathfrak{X}_{0}\cup\Phi_{flat}\mathfrak{X}_{0}$. Then, the following conditions are equivalent for a $kG$-module $M$:
(i) $M\in{\tt GProj}(kG)^{\perp}$,
(ii) $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$.
Proof. The implication (i)$\rightarrow$(ii) follows from Corollary 2.2(i), whereas the implication (ii)$\rightarrow$(i) follows from Theorem 3.1. $\Box$
We shall prove in the next Section (Proposition 4.3) that $\mathfrak{X}$ contains all groups $G$, for which $kG$ is weakly Gorenstein regular. In particular, $\mathfrak{X}$ contains all groups $G$, for which $kG$ is Gorenstein regular. Hence, the following result generalizes [29, Theorem 2.28].
Proposition 3.3.
Let $\mathfrak{X}_{0}$ be a subgroup-closed subclass of $\mathfrak{X}$, such that for any $\mathfrak{X}_{0}$-group $H$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kH),{\tt GProj}(kH)^{\perp}\right)$ is cogenerated by a set ${\tt S}_{H}$. Consider the closure $\overline{\mathfrak{X}_{0}}$ of $\mathfrak{X}_{0}$ under the operation ${\scriptstyle{{\bf LH}}}\cup\Phi\cup\Phi_{flat}$, as in §$\S$1.IV. Then, for any $\overline{\mathfrak{X}_{0}}$-group $G$ the Gorenstein projective cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is cogenerated by the set ${\tt T}_{G}=\bigcup\left\{\mbox{ind}_{H}^{G}{\tt S}_{H}:H\subseteq G,H\in\mathfrak{X}_{0}\right\}$.
Proof. Groups contained in $\overline{\mathfrak{X}_{0}}$ admit a hierarchical description, as explained in §$\S$1.IV. More precisely, we define the classes $\mathfrak{X}_{0,\alpha}$ for any ordinal number $\alpha$, as follows:
(a) $\mathfrak{X}_{0,0}=\mathfrak{X}_{0}$,
(b) $\mathfrak{X}_{0,\alpha+1}={\scriptstyle{{\bf LH}}}\mathfrak{X}_{0,\alpha}\cup\Phi\mathfrak{X}_{0,\alpha}\cup\Phi_{flat}\mathfrak{X}_{0,\alpha}$ for any ordinal $\alpha$ and
(c) $\mathfrak{X}_{0,\alpha}=\bigcup_{\beta<\alpha}\mathfrak{X}_{0,\beta}$ for any limit ordinal $\alpha$.
Then, $\overline{\mathfrak{X}_{0}}$ is the class consisting of those groups $G$, for which $G\in\mathfrak{X}_{0,\alpha}$ for some ordinal $\alpha$. We may therefore prove the claim in the statement of the Proposition using induction on the least ordinal $\alpha$, for which $G\in\mathfrak{X}_{0.\alpha}$. In view of Theorem 2.1(ii), the class $\overline{\mathfrak{X}_{0}}$ is a subclass of $\mathfrak{X}$; in particular, this is also the case for the class $\mathfrak{X}_{0,\alpha}$ for all $\alpha$.
First of all, we assume that $G\in\mathfrak{X}_{0,0}=\mathfrak{X}_{0}$. Since $\mbox{ind}_{H}^{G}{\tt S}_{H}\subseteq\mbox{ind}_{H}^{G}{\tt GProj}(kH)\subseteq{\tt GProj}(kG)$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$, we have ${\tt T}_{G}\subseteq{\tt GProj}(kG)$. As $G$ is an $\mathfrak{X}_{0}$-subgroup of itself, it is clear that ${\tt S}_{G}\subseteq{\tt T}_{G}$. The inclusions ${\tt S}_{G}\subseteq{\tt T}_{G}\subseteq{\tt GProj}(kG)$ imply that ${\tt GProj}(kG)^{\perp}\subseteq{\tt T}_{G}^{\perp}\subseteq{\tt S}_{G}^{\perp}$. Since ${\tt GProj}(kG)^{\perp}={\tt S}_{G}^{\perp}$, we conclude that ${\tt GProj}(kG)^{\perp}={\tt T}_{G}^{\perp}$, as needed.
For the inductive step of the proof, we assume that $G$ is an $\mathfrak{X}_{0,\alpha+1}$-group and the result is known for all $\mathfrak{X}_{0,\alpha}$-groups. For any $\mathfrak{X}_{0,\alpha}$-group $K$ our inductive assumption implies that ${\tt T}_{K}$ cogenerates the Gorenstein projective cotorsion pair over $kK$, i.e. we have ${\tt T}_{K}^{\perp}={\tt GProj}(kK)^{\perp}$. Since $\mathfrak{X}_{0,\alpha}$ is a subgroup-closed subclass of $\mathfrak{X}$ and
$G\in\mathfrak{X}_{0,\alpha+1}={\scriptstyle{{\bf LH}}}\mathfrak{X}_{0,\alpha}\cup\Phi\mathfrak{X}_{0,\alpha}\cup\Phi_{flat}\mathfrak{X}_{0,\alpha},$
we may apply Corollary 3.2 and conclude that the right orthogonal class ${\tt GProj}(kG)^{\perp}$ consists of those $kG$-modules $M$, for which $\mbox{res}_{K}^{G}M\in{\tt GProj}(kK)^{\perp}={\tt T}_{K}^{\perp}$ for any $\mathfrak{X}_{0,\alpha}$-subgroup $K\subseteq G$. In view of the induction-restriction (Eckmann-Shapiro) isomorphism, we have
$\mbox{res}_{K}^{G}M\in{\tt T}_{K}^{\perp}\Longleftrightarrow M\in\left(\mbox{ind}_{K}^{G}{\tt T}_{K}\right)\!^{\perp}.$
It follows that ${\tt GProj}(kG)^{\perp}=\bigcap\left\{\left(\mbox{ind}_{K}^{G}{\tt T}_{K}\right)\!^{\perp}:K\subseteq G,K\in\mathfrak{X}_{0,\alpha}\right\}$ and hence the Gorenstein projective cotorsion pair over $kG$ is cogenerated by $\,\bigcup\left\{\mbox{ind}_{K}^{G}{\tt T}_{K}:K\subseteq G,K\in\mathfrak{X}_{0,\alpha}\right\}$. On the other hand, for any $\mathfrak{X}_{0,\alpha}$-subgroup $K\subseteq G$, we have
$\mbox{ind}_{K}^{G}{\tt T}_{K}={\textstyle{\bigcup}}\left\{\mbox{ind}_{K}^{G}\mbox{ind}_{H}^{K}{\tt S}_{H}:H\subseteq K,H\in\mathfrak{X}_{0}\right\}\subseteq{\textstyle{\bigcup}}\left\{\mbox{ind}_{H}^{G}{\tt S}_{H}:H\subseteq G,H\in\mathfrak{X}_{0}\right\}.$
It follows that
${\textstyle{\bigcup}}\left\{\mbox{ind}_{K}^{G}{\tt T}_{K}:K\subseteq G,K\in\mathfrak{X}_{0,\alpha}\right\}\subseteq\,{\textstyle{\bigcup}}\left\{\mbox{ind}_{H}^{G}{\tt S}_{H}:H\subseteq G,H\in\mathfrak{X}_{0}\right\}={\tt T}_{G}.$
The latter inclusion is actually an equality, since $\mathfrak{X}_{0}\subseteq\mathfrak{X}_{0,\alpha}$ and ${\tt S}_{H}\subseteq{\tt T}_{H}$ for any $\mathfrak{X}_{0}$-subgroup $H\subseteq G$. Hence, the Gorenstein projective cotorsion pair over $kG$ is cogenerated by ${\tt T}_{G}$. $\Box$
There are results analogous to Theorem 3.1 and Corollary 3.2, regarding the left orthogonal to Gorenstein injective modules; these results will be used in the next section. For completeness, we shall provide the necessary arguments here. Consider the class $\mathfrak{Y}$ defined in Section 2; in view of Theorem 2.1(iii), $\mathfrak{Y}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{inj}$.
Theorem 3.4.
Let $\mathfrak{Y}_{0}$ be a subclass of $\mathfrak{Y}$. We consider a group $G$ and a $kG$-module $M$, such that $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ for any $\mathfrak{Y}_{0}$-subgroup $H\subseteq G$.
(i) We have $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ for any ${\scriptstyle{{\bf H}}}\mathfrak{Y}_{0}$-subgroup $H\subseteq G$.
(ii) If $\mathfrak{Y}_{0}$ is subgroup-closed, then $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ for any ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-subgroup $H\subseteq G$.
(iii) We have $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ for any $\Phi_{inj}\mathfrak{Y}_{0}$-subgroup $H\subseteq G$.
Proof. (i) We use induction on the ordinal $\alpha$, for which $H\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{Y}_{0}$. The base case where $\alpha=0$ is precisely our hypothesis on $M$. For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{Y}_{0}$-subgroups of $G$ and all $\beta<\alpha$. We consider an ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{Y}_{0}$-subgroup $H\subseteq G$ and let $N\in{\tt GInj}(kH)$. Then, as noted in §$\S$1.IV, there exists an exact sequence of $kH$-modules
$0\longrightarrow N\longrightarrow N^{0}\longrightarrow N^{1}\longrightarrow\ldots\longrightarrow N^{d}\longrightarrow 0,$
where $d$ is the dimension of the $H$-CW-complex witnessing that $H\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{Y}_{0}$ and each $N^{i}$ is a direct product of $kH$-modules of the form $\mbox{coind}_{F}^{H}\mbox{res}_{F}^{H}N$, with $F$ an ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{Y}_{0}$-subgroup of $H$ for some $\beta<\alpha$. Since ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{Y}_{0}\subseteq{\scriptstyle{{\bf H}}}_{\beta}\mathfrak{Y}\subseteq{\scriptstyle{{\bf H}}}\mathfrak{Y}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{Y}$, we may invoke property (4’) in Section 2 and conclude that restriction maps ${\tt GInj}(kH)$ into ${\tt GInj}(kF)$; in particular, it follows that $\mbox{res}_{F}^{H}N\in{\tt GInj}(kF)$. Our induction hypothesis implies that $\mbox{res}_{F}^{G}M\in{{}^{\perp}}{\tt GInj}(kF)$ and hence
$\mbox{Ext}^{n}_{kH}\!\left(\mbox{res}_{H}^{G}M,\mbox{coind}_{F}^{H}\mbox{res}_{F}^{H}N\right)=\mbox{Ext}^{n}_{kF}\!\left(\mbox{res}_{F}^{G}M,\mbox{res}_{F}^{H}N\right)=0$
for such a subgroup $F$ and all $n>0$. It follows that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N^{i}\right)=0$ for all $i$ and all $n>0$. Then, using dimension shifting, we conclude that $\mbox{Ext}_{kH}^{n}\!\left(\mbox{res}_{H}^{G}M,N\right)=0$ for all $n>d$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the kH-module $\mbox{res}_{H}^{G}M$ is actually contained in ${}^{\perp}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the $d$-th cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
$\mbox{Ext}_{kH}^{1}\!\left(\mbox{res}_{H}^{G}M,N\right)=\mbox{Ext}_{kH}^{d+1}\!\left(\mbox{res}_{H}^{G}M,N^{\prime}\right)=0.$
This completes the inductive step of the proof, taking care of any ${\scriptstyle{{\bf H}}}\mathfrak{Y}_{0}$-subgroup $H$ of $G$.
(ii) We shall now assume that $\mathfrak{Y}_{0}$ is subgroup-closed and consider an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-subgroup $H$ of $G$. We proceed by induction on the cardinality $\kappa$ of $H$. If $\kappa$ is countable, then the ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group $H$ is actually contained in ${\scriptstyle{{\bf H}}}\mathfrak{Y}_{0}$ and we are done by (i) above. If $\kappa$ is uncountable, then we may express $H$ as a continuous ascending union of subgroups $(H_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Since the class ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$ is subgroup-closed, $H_{\lambda}$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}$-group; hence, our induction hypothesis implies that $\mbox{res}^{G}_{H_{\lambda}}M\in{{}^{\perp}}{\tt GInj}(kH_{\lambda})$ for all $\lambda$. We now fix $N\in{\tt GInj}(kH)$ and note that $\mbox{res}^{H}_{H_{\lambda}}N\in{\tt GInj}(kH_{\lambda})$ for all $\lambda$; this follows since $H_{\lambda}\in{\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{Y}$, in view of property (4’) in Section 2. Letting $M_{\lambda}=\mbox{ind}_{H_{\lambda}}^{H}\mbox{res}^{G}_{H_{\lambda}}M$, we note that
$\mbox{Ext}^{n}_{kH}\!\left(M_{\lambda},N\right)=\mbox{Ext}^{n}_{kH_{\lambda}}\!\left(\mbox{res}^{G}_{H_{\lambda}}M,\mbox{res}_{H_{\lambda}}^{H}N\right)=0$
for all $\lambda$ and $n>0$. The ascending family of subgroups $(H_{\lambda})_{\lambda<\kappa}$ induces a continuous direct system of $kH$-modules $(M_{\lambda})_{\lambda<\kappa}$ with surjective structure maps, whose colimit is $\mbox{res}_{H}^{G}M$. The short exact sequence of $kH$-modules
$0\longrightarrow T_{\lambda}\longrightarrow M_{0}\longrightarrow M_{\lambda}\longrightarrow 0,$
where $T_{\lambda}$ is the kernel of the structure map $M_{0}\longrightarrow M_{\lambda}$, is the $\lambda$-th term of a continuous direct system of short exact sequences. The colimit of the latter direct system of short exact sequences is the short exact sequence of $kH$-modules
(2) $0\longrightarrow T\longrightarrow M_{0}\longrightarrow\mbox{res}_{H}^{G}M\longrightarrow 0$
Here, $T$ is equal to the continuous ascending union of its submodules $(T_{\lambda})_{\lambda<\kappa}$. Since $T_{\lambda+1}/T_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kH}\!\left(T_{\lambda+1}/T_{\lambda},N\right)=0$ for all $\lambda$ and $n>0$. Therefore, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kH}(T,N)=0$ for all $n>0$. Then, the short exact sequence (2) above implies that $\mbox{Ext}^{n}_{kH}\!\left(\mbox{res}_{H}^{G}M,N\right)\!=0$ for all $n>1$. Since this holds for any $N\in{\tt GInj}(kH)$, we can easily conclude that the $kH$-module $\mbox{res}_{H}^{G}M$ is actually contained in ${{}^{\perp}}{\tt GInj}(kH)$. Indeed, the Gorenstein injective $kH$-module $N$ is the first cosyzygy in an injective resolution of another Gorenstein injective $kH$-module $N^{\prime}$ (this is property (2’) in Section 2) and hence
$\mbox{Ext}_{kH}^{1}\!\left(\mbox{res}_{H}^{G}M,N\right)\!=\mbox{Ext}_{kH}^{2}\!\left(\mbox{res}_{H}^{G}M,N^{\prime}\right)\!=0.$
This completes the proof of (ii).
(iii) We fix a $\Phi_{inj}\mathfrak{Y}_{0}$-subgroup $H\subseteq G$. Since the cotorsion pair $\left({{}^{\perp}}{\tt GInj}(kH),{\tt GInj}(kH)\right)$ is complete (property (5’) in Section 2), there exists a short exact sequence of $kH$-modules
$0\longrightarrow N\longrightarrow L\longrightarrow\mbox{res}_{H}^{G}M\longrightarrow 0,$
where $L\in{{}^{\perp}}{\tt GInj}(kH)$ and $N\in{\tt GInj}(kH)$. Then, for any $\mathfrak{Y}_{0}$-subgroup $F\subseteq H$ we obtain the exact sequence of restricted $kF$-modules
$0\longrightarrow\mbox{res}_{F}^{H}N\longrightarrow\mbox{res}_{F}^{H}L\longrightarrow\mbox{res}_{F}^{G}M\longrightarrow 0$
Corollary 2.2(ii) implies that $\mbox{res}_{F}^{H}L\in{{}^{\perp}}{\tt GInj}(kF)$, whereas property (4’) of Section 2 shows that $\mbox{res}_{F}^{H}N\in{\tt GInj}(kF)$. Since $\mbox{res}_{F}^{G}M\in{{}^{\perp}}{\tt GInj}(kF)$, in view of our assumption on $M$, the thickness of the class ${{}^{\perp}}{\tt GInj}(kF)$ implies that $\mbox{res}_{F}^{H}N$ is also contained in ${{}^{\perp}}{\tt GInj}(kF)$. Since ${\tt GInj}(kF)\cap{{}^{\perp}}{\tt GInj}(kF)={\tt Inj}(kF)$, it follows that $\mbox{res}_{F}^{H}N$ is an injective $kF$-module. This is the case for any $\mathfrak{Y}_{0}$-subgroup $F$ of the $\Phi_{inj}\mathfrak{Y}_{0}$-group $H$ and hence $\mbox{id}_{kH}N<\infty$. It follows that $N\in{{}^{\perp}}{\tt GInj}(kH)$ and hence the thickness of ${{}^{\perp}}{\tt GInj}(kH)$ implies that $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$, as needed. $\Box$
Corollary 3.5.
Let $\mathfrak{Y}_{0}$ be a subgroup-closed subclass of $\mathfrak{Y}$ and consider a group $G$, which is contained in ${\scriptstyle{{\bf LH}}}\mathfrak{Y}_{0}\cup\Phi_{inj}\mathfrak{Y}_{0}$. Then, the following are equivalent for a $kG$-module $M$:
(i) $M\in{{}^{\perp}}{\tt GInj}(kG)$,
(ii) $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ for any $\mathfrak{Y}_{0}$-subgroup $H\subseteq G$.
Proof. The implication (i)$\rightarrow$(ii) follows from Corollary 2.2(ii), whereas the implication (ii)$\rightarrow$(i) follows from Theorem 3.4. $\Box$
4. Virtually Gorenstein group algebras
We fix a commutative ring $k$ and consider groups $G$ for which the group algebra $kG$ is virtually Gorenstein, i.e. for which ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. This homological condition on $kG$ has certain interesting consequences regarding the (homotopy) properties of Gorenstein modules. As an indication of these properties, we note that if $kG$ is virtually Gorenstein, then: [^2] Properties (a), (b), (c) and (d) are actually valid over any (left) virtually Gorenstein ring $R$.
(a) (cf. [37, Proposition 3.16]) There is an equality ${\tt GProj}(kG)={\tt PGF}(kG)$; hence, [35, Theorem 4.9] implies that the cotorsion pair $\left({\tt GProj}(kG),{\tt GProj}(kG)^{\perp}\right)$ is complete.
(b) (cf. [11, Proposition 2.6]) Let $M$ be a $kG$-module. Any resolution of $M$ which is obtained by iterating special Gorenstein projective precovers remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$. Dually, any coresolution of $M$ which is obtained by iterating special Gorenstein injective preenvelopes remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(c) (cf. [11, Proposition 2.2]) An acyclic complex of $kG$-modules remains acyclic by applying the functor $\mbox{Hom}_{kG}(\_\!\_,N)$ for any Gorenstein injective $kG$-module $N$ if and only if it remains acyclic by applying the functor $\mbox{Hom}_{kG}(L,\_\!\_)$ for any Gorenstein projective $kG$-module $L$.
(d) (cf. [37, Example 3.18(2)]) Let $\mathcal{C}$ be the class of those acyclic complexes of $kG$-modules that satisfy the equivalent conditions described in (c) above. Then, the relative derived category, which is obtained from the homotopy category of $kG$ by inverting those chain maps whose mapping cone is a $\mathcal{C}$-complex, is triangulated equivalent to certain homotopy categories of Gorenstein projective (resp. Gorenstein injective) $kG$-modules.
We consider the classes $\mathfrak{X}$ and $\mathfrak{Y}$ defined in Section 2 and let $\mathfrak{Z}=\mathfrak{Z}(k)$ be the class consisting of those groups $G\in\mathfrak{X}\cap\mathfrak{Y}$, for which $kG$ is virtually Gorenstein.
Lemma 4.1.
The class $\mathfrak{Z}$ is subgroup-closed.
Proof. Let $G$ be a $\mathfrak{Z}$-group and consider a subgroup $H\subseteq G$. Then, $G\in\mathfrak{X}\cap\mathfrak{Y}$ and the group algebra $kG$ is virtually Gorenstein. Since $\mathfrak{X}$ and $\mathfrak{Y}$ are subgroup-closed classes (cf. Theorem 2.1), it follows that $H\in\mathfrak{X}\cap\mathfrak{Y}$. It remains to show that $kH$ is virtually Gorenstein.
Consider a $kH$-module $M\in{\tt GProj}(kH)^{\perp}$. Since $G\in\mathfrak{Z}\subseteq\mathfrak{X}$ and $\mathfrak{X}$ is subgroup-closed, we conclude that $H$ is an $\mathfrak{X}$-group. It follows that restriction maps ${\tt GProj}(kG)$ into ${\tt GProj}(kH)$ (this is precisely property (4) in Section 2), so that the coinduced $kG$-module $\mbox{coind}_{H}^{G}M$ is contained in ${\tt GProj}(kG)^{\perp}$, i.e. in ${{}^{\perp}}{\tt GInj}(kG)$. Invoking Corollary 2.2(ii), we may now conclude that $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}M$, it follows that $M\in{{}^{\perp}}{\tt GInj}(kH)$. We have therefore shown that ${\tt GProj}(kH)^{\perp}\subseteq{{}^{\perp}}{\tt GInj}(kH)$.
Conversely, assume that $M\in{{}^{\perp}}{\tt GInj}(kH)$. Since $G\in\mathfrak{Z}\subseteq\mathfrak{Y}$ and $\mathfrak{Y}$ is subgroup-closed, we may conclude that $H$ is a $\mathfrak{Y}$-group. Hence, restriction maps ${\tt GInj}(kG)$ into ${\tt GInj}(kH)$ (this is precisely property (4’) in Section 2), so that the induced $kG$-module $\mbox{ind}_{H}^{G}M$ is contained in ${{}^{\perp}}{\tt GInj}(kG)$, i.e. in ${\tt GProj}(kG)^{\perp}$. Corollary 2.2(i) then implies that $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$. Since $M$ is a direct summand of $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}M$, it follows that $M\in{\tt GProj}(kH)^{\perp}$. We have thus shown that ${{}^{\perp}}{\tt GInj}(kH)\subseteq{\tt GProj}(kH)^{\perp}$. $\Box$
The following result will provide us with an abundance of examples of virtually Gorenstein group algebras.
Theorem 4.2.
The class $\mathfrak{Z}$ is closed under the operations (i) ${\scriptstyle{{\bf LH}}}$ and (ii) $\Phi_{inj}$.
Proof. (i) We have to show that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}=\mathfrak{Z}$, i.e. that ${\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq\mathfrak{Z}$. To that end, we consider an ${\scriptstyle{{\bf LH}}}\mathfrak{Z}$-group $G$ and aim at proving that $G\in\mathfrak{Z}$. First of all, we note that
$G\in{\scriptstyle{{\bf LH}}}\mathfrak{Z}\subseteq{\scriptstyle{{\bf LH}}}(\mathfrak{X}\cap\mathfrak{Y})\subseteq{\scriptstyle{{\bf LH}}}\mathfrak{X}\cap{\scriptstyle{{\bf LH}}}\mathfrak{Y}=\mathfrak{X}\cap\mathfrak{Y}.$
For any $\mathfrak{Z}$-subgroup $H\subseteq G$ we denote by $\left(\mbox{res}_{H}^{G}\right)^{-1}{\tt GProj}(kH)^{\perp}$ the class of those $kG$-modules $M$ for which $\mbox{res}_{H}^{G}M\in{\tt GProj}(kH)^{\perp}$ and apply Corollary 3.2 (with $\mathfrak{X}_{0}=\mathfrak{Z}\subseteq\mathfrak{X}$ therein), in order to conclude that
${\tt GProj}(kG)^{\perp}={\textstyle{\bigcap}}\left\{\left(\mbox{res}_{H}^{G}\right)^{-1}{\tt GProj}(kH)^{\perp}:H\subseteq G,H\in\mathfrak{Z}\right\}.$
Analogously, for any $\mathfrak{Z}$-subgroup $H\subseteq G$ we denote by $\left(\mbox{res}_{H}^{G}\right)^{-1}{{}^{\perp}}{\tt GInj}(kH)$ the class of those $kG$-modules $M$ for which $\mbox{res}_{H}^{G}M\in{{}^{\perp}}{\tt GInj}(kH)$ and apply Corollary 3.5 (with $\mathfrak{Y}_{0}=\mathfrak{Z}\subseteq\mathfrak{Y}$ therein), in order to conclude that
${{}^{\perp}}{\tt GInj}(kG)={\textstyle{\bigcap}}\left\{\left(\mbox{res}_{H}^{G}\right)^{-1}{{}^{\perp}}{\tt GInj}(kH):H\subseteq G,H\in\mathfrak{Z}\right\}.$
Since ${\tt GProj}(kH)^{\perp}={{}^{\perp}}{\tt GInj}(kH)$ for any $\mathfrak{Z}$-subgroup $H\subseteq G$, the above equalities show that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$. The group algebra $kG$ is therefore virtually Gorenstein and hence $G\in\mathfrak{Z}$.
(ii) In order to show that $\Phi_{inj}\mathfrak{Z}=\mathfrak{Z}$, i.e. that $\Phi_{inj}\mathfrak{Z}\subseteq\mathfrak{Z}$, we consider a $\Phi_{inj}\mathfrak{Z}$-group $G$ and aim at proving that $G\in\mathfrak{Z}$. We have
$G\in\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{inj}(\mathfrak{X}\cap\mathfrak{Y})\subseteq\Phi_{inj}\mathfrak{X}\cap\Phi_{inj}\mathfrak{Y}\subseteq\Phi_{flat}\mathfrak{X}\cap\Phi_{inj}\mathfrak{Y}=\mathfrak{X}\cap\mathfrak{Y},$
where the third inclusion follows since $\Phi_{inj}\mathfrak{X}\subseteq\Phi_{flat}\mathfrak{X}$; cf. [22, Lemma 1.5]. The group $G$ is contained in $\Phi_{inj}\mathfrak{Z}\subseteq\Phi_{flat}\mathfrak{Z}$ and hence we can apply Corollaries 3.2 and 3.5 in exactly the same way as in (i) above, in order to show that ${\tt GProj}(kG)^{\perp}={{}^{\perp}}{\tt GInj}(kG)$, i.e. that the group algebra $kG$ is virtually Gorenstein. This completes the proof that $G\in\mathfrak{Z}$. $\Box$
Let $\mathfrak{W}=\mathfrak{W}(k)$ be the class consisting of those groups $G$, for which $kG$ is weakly Gorenstein regular; in other words, $G\in\mathfrak{W}$ if the invariant $\mbox{sfli}\,kG$ is finite. This class contains all groups $G$, for which the group algebra $kG$ is Gorenstein regular. We note that $\mathfrak{W}$ is subgroup-closed; indeed, for any subgroup $H$ of a group $G$ there is an inequality $\mbox{sfli}\,kH\leq\mbox{sfli}\,kG$. Hence, $\mathfrak{W}$ is non-empty if and only if it contains the trivial group, i.e. if and only if the coefficient ring $k$ is itself weakly Gorenstein regular. In that case, $\mathfrak{W}$ is known to contain the class $\mathfrak{F}$ of finite groups; see, for example, [33].
Proposition 4.3.
The class $\mathfrak{W}$ is a subclass of $\mathfrak{Z}$.
Proof. We note the following general facts, regarding a group $G$: If $F_{*}$ is an acyclic complex of flat $kG$-modules, then the complex of abelian groups $M\otimes_{kG}F_{*}$ is acyclic for any $kG$-module $M$ of finite flat dimension; this follows by induction on $\mbox{fd}_{kG}M$. Similarly, if $I^{*}$ is an acyclic complex of injective $kG$-modules and $M$ is a $kG$-module of finite flat dimension, then the complex of abelian groups $\mbox{Hom}_{kG}(M,I^{*})$ is acyclic; cf. [34, Corollary 5.9].
In order to prove that $\mathfrak{W}\subseteq\mathfrak{Z}$, we have to show that any $\mathfrak{W}$-group $G$ has the following three properties:
(i) $G\in\mathfrak{X}=\mathfrak{X}^{\prime}$, i.e. ${\tt GFlat}(kG)=\mathcal{F}(kG)$,
(ii) $G\in\mathfrak{Y}$, i.e. ${\tt GInj}(kG)=\mathcal{I}(kG)$ and
(iii) $kG$ is virtually Gorenstein.
Since $\mbox{sfli}\,kG$ is finite, all injective $kG$-modules have finite flat dimension. Hence, properties (i) and (ii) follow from the two general facts mentioned above; see also [39, Theorem 2.3 and Theorem 2.6]. As far as property (iii) is concerned, this is precisely [18, Theorem A]; see also [39, Lemma 3.2]. $\Box$
Let $\overline{\mathfrak{W}}$ be the smallest class of groups which contains $\mathfrak{W}$ and is ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Groups in $\overline{\mathfrak{W}}$ admit a hierarchical description, as explained in §$\S 1$.IV. The following Corollary implies Theorem A stated in the Introduction; it is itself an immediate consequence of Theorem 4.2 and Proposition 4.3.
Corollary 4.4.
The class $\overline{\mathfrak{W}}$ is contained in $\mathfrak{Z}$; in particular, $kG$ is virtually Gorenstein for any $\overline{\mathfrak{W}}$-group $G$.
5. Groups satisfying Moore’s conjecture
Motivated by Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [10, Chapter VIII, §$\S$3], Moore has conjectured that a module $M$ over the group algebra of a group $G$ is projective, whenever the restriction of $M$ to a subgroup $H\subseteq G$ of finite index that satisfies an additional group-theoretic condition is projective. More precisely, we say that the triple $(G,H,k)$, where $G$ is a group and $H\subseteq G$ is a subgroup of finite index, satisfies Moore’s condition if for any element $g\in G\setminus H$ one of the following two conditions hold:
(i) the intersection $\langle g\rangle\cap H$ is non-trivial or
(ii) $g$ has finite order, which is invertible in $k$.
Following [1], we say that $G$ satisfies Moore’s conjecture over $k$ if for any subgroup $H\subseteq G$ of finite index, such that the triple $(G,H,k)$ satisfies Moore’s condition, and any $kG$-module $M$, for which $\mbox{res}_{H}^{G}M\in{\tt Proj}(kH)$, we actually have $M\in{\tt Proj}(kG)$. We denote by $\mathfrak{M}=\mathfrak{M}(k)$ the class consisting of those groups that satisfy Moore’s conjecture over $k$. Replacing projective modules by flat (resp. injective) modules in the above discussion, we obtain versions of Moore’s conjecture for flat (resp. injective) modules; cf. [2]. We shall denote by $\mathfrak{M}_{flat}=\mathfrak{M}_{flat}(k)$ and $\mathfrak{M}_{inj}=\mathfrak{M}_{inj}(k)$ the corresponding classes of groups. In other words, $\mathfrak{M}_{flat}$ (resp. $\mathfrak{M}_{inj}$) is the class consisting of those groups that satisfy the flat (resp. injective) version of Moore’s conjecture over $k$.
Proposition 5.1.
There are inclusions $\mathfrak{M}_{inj}\subseteq\mathfrak{M}_{flat}\subseteq\mathfrak{M}$.
Proof. The first inclusion $\mathfrak{M}_{inj}\subseteq\mathfrak{M}_{flat}$ is a consequence of Lambek’s flatness criterion [31], as noted in [2, Proposition 2.6]. The inclusion $\mathfrak{M}_{flat}\subseteq\mathfrak{M}$ follows from the Benson-Goodearl theorem [8, Corollary 4.8], as in the proof of [2, Proposition 2.7]. $\Box$
The following simple (and well-known) result will be helpful in the proof of Theorems 5.3-5.5 below. We include a proof for the sake of completeness.
Proposition 5.2.
Let $G$ be a group, $H\subseteq G$ a subgroup of finite index and $M$ a $kG$-module.
(i) If $\mbox{pd}_{kG}M<\infty$ and $\mbox{res}_{H}^{G}M\in{\tt Proj}(kH)$, then $M$ is projective.
(ii) If $\mbox{id}_{kG}M<\infty$ and $\mbox{res}_{H}^{G}M\in{\tt Inj}(kH)$, then $M$ is injective.
(iii) If $\mbox{fd}_{kG}M<\infty$ and $\mbox{res}_{H}^{G}M\in{\tt Flat}(kH)$, then $M$ is flat.
(iv) If $\mbox{fd}_{kG}M<\infty$ and $\mbox{res}_{H}^{G}M\in{\tt Proj}(kH)$, then $M$ is projective.
Proof. (i) Assume that $n=\mbox{pd}_{kG}M\geq 1$ and consider an exact sequence of $kG$-modules
$0\longrightarrow P_{n}\stackrel{{\scriptstyle f}}{{\longrightarrow}}P_{n-1}\longrightarrow\ldots\longrightarrow P_{0}\longrightarrow M\longrightarrow 0,$
where $P_{0},\ldots,P_{n-1},P_{n}\in{\tt Proj}(kG)$. We denote this exact sequence by $X_{*}$. Since $f$ is not a split monomorphism, the complex of abelian groups $\mbox{Hom}_{kG}(X_{*},P_{n})$ is not acyclic. We shall reach a contradiction, by showing that $\mbox{Hom}_{kG}(X_{*},P)$ is acyclic for any projective $kG$-module $P$. In fact, it suffices to consider the case where $P$ is $kG$-free. Then, $P=\mbox{ind}_{H}^{G}Q=\mbox{coind}_{H}^{G}Q$ for a suitable (free) $kH$-module $Q$ and hence
$\mbox{Hom}_{kG}(X_{*},P)=\mbox{Hom}_{kG}\!\left(X_{*},\mbox{coind}_{H}^{G}Q\right)\!=\mbox{Hom}_{kH}\!\left(\mbox{res}_{H}^{G}X_{*},Q\right)\!.$
In view of our assumption on $\mbox{res}_{H}^{G}M$, the restricted complex $\mbox{res}_{H}^{G}X_{*}$ is contractible (being a right-bounded acyclic complex of projective $kH$-modules). Therefore, the complex of abelian groups $\mbox{Hom}_{kG}(X_{*},P)=\mbox{Hom}_{kH}\!\left(\mbox{res}_{H}^{G}X_{*},Q\right)$ is contractible as well; in particular, it is acyclic.
(ii) Assume that $n=\mbox{id}_{kG}M\geq 1$ and consider an exact sequence of $kG$-modules
$0\longrightarrow M\longrightarrow I^{0}\longrightarrow\ldots\longrightarrow I^{n-1}\stackrel{{\scriptstyle g}}{{\longrightarrow}}I^{n}\longrightarrow 0,$
where $I^{0},\ldots,I^{n-1},I^{n}\in{\tt Inj}(kG)$. We denote this exact sequence by $Y^{*}$. Since $g$ is not a split epimorphism, the complex of abelian groups $\mbox{Hom}_{kG}(I^{n},Y^{*})$ is not acyclic. We shall reach a contradiction, by showing that $\mbox{Hom}_{kG}(I,Y^{*})$ is acyclic for any injective $kG$-module $I$. In fact, it suffices to consider the case where $I$ is a coinduced module of the form $\mbox{coind}_{1}^{G}A$, for a divisible abelian group $A$. Then, $I=\mbox{coind}_{H}^{G}J=\mbox{ind}_{H}^{G}J$ for a suitable (injective) $kH$-module $J$ and hence
$\mbox{Hom}_{kG}(I,Y^{*})=\mbox{Hom}_{kG}\!\left(\mbox{ind}_{H}^{G}J,Y^{*}\right)\!=\mbox{Hom}_{kH}\!\left(J,\mbox{res}_{H}^{G}Y^{*}\right)\!.$
In view of our assumption on $\mbox{res}_{H}^{G}M$, the restricted complex $\mbox{res}_{H}^{G}Y^{*}$ is contractible (being a left-bounded acyclic complex of injective $kH$-modules). It follows that the complex of abelian groups $\mbox{Hom}_{kG}(I,Y^{*})=\mbox{Hom}_{kH}\!\left(J,\mbox{res}_{H}^{G}Y^{*}\right)$ is contractible as well; in particular, it is acyclic.
(iii) This follows from assertion (ii) above, in view of Lambek’s flatness criterion [31]. Indeed, the Pontryagin dual $DM$ of $M$ has injective dimension $\mbox{id}_{kG}DM=\mbox{fd}_{kG}M<\infty$, whereas the restricted $kH$-module $\mbox{res}_{H}^{G}DM=D\mbox{res}_{H}^{G}M$ is injective. We may then invoke (ii) and conclude that $DM\in{\tt Inj}(kG)$; hence, $M\in{\tt Flat}(kG)$.
(iv) Assume that $\mbox{fd}_{kG}M=n$ and let $N=\Omega^{n}M$ be the $n$-th syzygy in a projective resolution of $M$. Then, the $kG$-module $N$ is flat and its restriction $\mbox{res}_{H}^{G}N=\Omega^{n}\mbox{res}_{H}^{G}M$ is $kH$-projective. Invoking the Benson-Goodearl theorem [8, Corollary 4.8], we conclude that $N\in{\tt Proj}(kG)$ and hence $\mbox{pd}_{kG}M\leq n<\infty$. Then, the projectivity of $M$ is a consequence of (i) above. $\Box$
The next results are based upon and, at the same time, complement previous work on Moore’s conjecture by Aljadeff, Meir, Bahlekeh and Salarian; cf. [1], [2].
Theorem 5.3.
Let $G$ be a group and $H\subseteq G$ a subgroup of finite index, such that the triple $(G,H,k)$ satisfies Moore’s condition. If $M$ is a $kG$-module with $\mbox{res}_{H}^{G}M\in{\tt Proj}(kH)$, then:
(i) $\mbox{res}_{S}^{G}M\in{\tt Proj}(kS)$ for any ${\scriptstyle{{\bf LH}}}\mathfrak{M}$-subgroup $S\subseteq G$,
(ii) $\mbox{res}_{S}^{G}M\in{\tt Proj}(kS)$ for any $\Phi\mathfrak{M}$-subgroup $S\subseteq G$ and
(iii) $\mbox{res}_{S}^{G}M\in{\tt Proj}(kS)$ for any $\Phi_{flat}\mathfrak{M}$-subgroup $S\subseteq G$.
Proof. For the subgroup $S$ under consideration, we let $N=\mbox{res}_{S}^{G}M$ and note that $\mbox{res}_{H\cap S}^{S}N=\mbox{res}_{H\cap S}^{G}M=\mbox{res}_{H\cap S}^{H}\mbox{res}_{H}^{G}M$. Then, being a restriction of the projective $kH$-module $\mbox{res}_{H}^{G}M$, the latter $k(H\cap S)$-module is projective as well. We also note that $H\cap S$ has finite index in $S$ and the triple $(S,H\cap S,k)$ satisfies Moore’s condition.
(i) We shall first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}$, following the argument in the proof of [1, Proposition 4.2]. If $\alpha=0$, then $S\in\mathfrak{M}$ and hence $S$ satisfies Moore’s conjecture over $k$. The $kS$-module $N$ is therefore projective (as its restriction to $H\cap S$ is projective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules
$0\longrightarrow N_{d}\longrightarrow\ldots\longrightarrow N_{1}\longrightarrow N_{0}\longrightarrow N\longrightarrow 0,$
where $d$ is the dimension of the $S$-CW-complex witnessing that $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}$ and each $N_{i}$ is a direct sum of $kS$-modules of the form $\mbox{ind}_{F}^{S}\mbox{res}_{F}^{S}N$, with $F$ an ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}$-subgroup of $S$ for some $\beta<\alpha$. Our inductive assumption implies that the $kF$-module $\mbox{res}_{F}^{S}N=\mbox{res}_{F}^{G}M$ is projective, so that $\mbox{ind}_{F}^{S}\mbox{res}_{F}^{S}N\in{\tt Proj}(kS)$ for all such subgroups $F$. Then, the $kS$-module $N_{i}$ is projective for all $i=0,1,\ldots,d$ and hence $\mbox{pd}_{kS}N\leq d$. Invoking Proposition 5.2(i), we conclude that the $kS$-module $N$ is actually projective.
We now assume that $S$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{M}$-subgroup of $G$. If $S^{\prime}\subseteq S$ is a finitely generated subgroup, then there exists an ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroup $S^{\prime\prime}\subseteq S$ with $S^{\prime}\subseteq S^{\prime\prime}$. Since the $kS^{\prime\prime}$-module $\mbox{res}_{S^{\prime\prime}}^{G}M$ is projective (in view of the above discussion, regarding ${\scriptstyle{{\bf H}}}\mathfrak{M}$-subgroups of $G$), this is also the case for its restriction $\mbox{res}_{S^{\prime}}^{G}M=\mbox{res}_{S^{\prime}}^{S}N$ to $S^{\prime}$; in particular, the $kS^{\prime}$-module $\mbox{res}_{S^{\prime}}^{S}N$ is flat. As $N$ is the filtered colimit of the $kS$-modules $\mbox{ind}_{S^{\prime}}^{S}\mbox{res}_{S^{\prime}}^{S}N\in{\tt Flat}(kS)$, where $S^{\prime}$ runs through all finitely generated subgroups of $S$, the $kS$-module $N$ is also flat. The projectivity of $N$ follows now from the Benson-Goodearl theorem [8, Corollary 4.8].
(ii) For any $\mathfrak{M}$-subgroup $F\subseteq S$, we consider the $kF$-module $\mbox{res}_{F}^{S}N=\mbox{res}_{F}^{G}M$ and note that its restriction $\mbox{res}_{H\cap F}^{F}N=\mbox{res}_{H\cap F}^{G}M=\mbox{res}_{H\cap S}^{H}\mbox{res}_{H}^{G}M$ to $H\cap F$ is projective (as a restriction of the projective $kH$-module $\mbox{res}_{H}^{G}M$). We also note that $(F,H\cap F,k)$ satisfies Moore’s condition, whereas the $\mathfrak{M}$-group $F$ satisfies Moore’s conjecture; it follows that $\mbox{res}_{F}^{S}N\in{\tt Proj}(kF)$. This holds for any $\mathfrak{M}$-subgroup $F$ of the $\Phi\mathfrak{M}$-group $S$ and hence $\mbox{pd}_{kS}N<\infty$. Then, Proposition 5.2(i) implies that $N\in{\tt Proj}(kS)$.
(iii) Working as in (ii) above, we conclude that the $kF$-module $\mbox{res}_{F}^{S}N$ is projective and hence flat for any $\mathfrak{M}$-subgroup $F$ of the $\Phi_{flat}\mathfrak{M}$-group $S$. Therefore, it follows that $\mbox{fd}_{kS}N<\infty$; an application of Proposition 5.2(iv) shows that $N$ is actually projective. $\Box$
Theorem 5.4.
Let $G$ be a group and $H\subseteq G$ a subgroup of finite index, such that the triple $(G,H,k)$ satisfies Moore’s condition. If $M$ is a $kG$-module with $\mbox{res}_{H}^{G}M\in{\tt Flat}(kH)$, then:
(i) $\mbox{res}_{S}^{G}M\in{\tt Flat}(kS)$ for any ${\scriptstyle{{\bf LH}}}\mathfrak{M}_{flat}$-subgroup $S\subseteq G$,
(ii) $\mbox{res}_{S}^{G}M\in{\tt Flat}(kS)$ for any $\Phi_{flat}\mathfrak{M}_{flat}$-subgroup $S\subseteq G$.
Proof. We can follow the arguments in the proof of assertions (i) and (ii) of Theorem 5.3, using Proposition 5.2(iii) instead of Proposition 5.2(i). We note that there is no need to appeal to the Benson-Goodearl theorem at the end of the proof of (i). $\Box$
Theorem 5.5.
Let $G$ be a group and $H\subseteq G$ a subgroup of finite index, such that the triple $(G,H,k)$ satisfies Moore’s condition. If $M$ is a $kG$-module with $\mbox{res}_{H}^{G}M\in{\tt Inj}(kH)$, then:
(i) $\mbox{res}_{S}^{G}M\in{\tt Inj}(kS)$ for any ${\scriptstyle{{\bf LH}}}\mathfrak{M}_{inj}$-subgroup $S\subseteq G$,
(ii) $\mbox{res}_{S}^{G}M\in{\tt Inj}(kS)$ for any $\Phi_{inj}\mathfrak{M}_{inj}$-subgroup $S\subseteq G$.
Proof. As we did in the proof of Theorem 5.3, for the subgroup $S$ under consideration, we let $N=\mbox{res}_{S}^{G}M$ and note that $\mbox{res}_{H\cap S}^{S}N=\mbox{res}_{H\cap S}^{G}M=\mbox{res}_{H\cap S}^{H}\mbox{res}_{H}^{G}M$. Then, being a restriction of the injective $kH$-module $\mbox{res}_{H}^{G}M$, the latter $k(H\cap S)$-module is injective as well. We also note that $H\cap S$ has finite index in $S$ and the triple $(S,H\cap S,k)$ satisfies Moore’s condition.
(i) We first consider the case where $S$ is an ${\scriptstyle{{\bf H}}}\mathfrak{M}_{inj}$-subgroup of $G$ and proceed by induction on the ordinal $\alpha$, for which $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}_{inj}$, following the argument in the proof of [2, Theorem 2.5]. If $\alpha=0$, then $S\in\mathfrak{M}_{inj}$ and hence $S$ satisfies the injective version of Moore’s conjecture. The $kS$-module $N$ is therefore injective (as its restriction to $H\cap S$ is injective). For the inductive step, assume that $\alpha>0$ and the result is true for all ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}_{inj}$-subgroups of $G$ and all $\beta<\alpha$. As we noted in §$\S$1.IV, there exists a short exact sequence of $kS$-modules
$0\longrightarrow N\longrightarrow N^{0}\longrightarrow N^{1}\longrightarrow\ldots\longrightarrow N^{d}\longrightarrow 0,$
where $d$ is the dimension of the $S$-CW-complex witnessing that $S\in{\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{M}_{inj}$ and each $N^{i}$ is a direct product of $kS$-modules of the form $\mbox{coind}_{F}^{S}\mbox{res}_{F}^{S}N$, with $F$ an ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{M}_{inj}$-subgroup of $S$ for some $\beta<\alpha$. Our inductive assumption implies that the $kF$-module $\mbox{res}_{F}^{S}N=\mbox{res}_{F}^{G}M$ is injective, so that $\mbox{coind}_{F}^{S}\mbox{res}_{F}^{S}N\in{\tt Inj}(kS)$ for all such subgroups $F$. Hence, $N^{i}$ is injective for all $i=0,1,\ldots,d$ and $\mbox{id}_{kS}N\leq d$. Finally, an application of Proposition 5.2(ii) implies that the $kS$-module $N$ is actually injective.
We now assume that $S$ is an ${\scriptstyle{{\bf LH}}}\mathfrak{M}_{inj}$-subgroup of $G$. For any finitely generated subgroup $S^{\prime}\subseteq S$, there exists an ${\scriptstyle{{\bf H}}}\mathfrak{M}_{inj}$-subgroup $S^{\prime\prime}\subseteq S$ with $S^{\prime}\subseteq S^{\prime\prime}$. Since the $kS^{\prime\prime}$-module $\mbox{res}_{S^{\prime\prime}}^{G}M$ is injective (in view of the above discussion, regarding ${\scriptstyle{{\bf H}}}\mathfrak{M}_{inj}$-subgroups of $G$), this is also the case for its restriction $\mbox{res}_{S^{\prime}}^{G}M=\mbox{res}_{S^{\prime}}^{S}N$ to $S^{\prime}$. This holds for any finitely generated subgroup $S^{\prime}\subseteq S$ and hence Lemma 5.6 below implies (by letting $H^{\prime}=H\cap S$ therein) that $N\in{\tt Inj}(kS)$.
(ii) We can follow the arguments in the proof of Theorem 5.3(ii), using Proposition 5.2(ii) instead of Proposition 5.2(i). $\Box$
The following Lemma was used in the proof of Theorem 5.5. We have separated its proof from the main body of the proof of that Theorem for expository reasons.
Lemma 5.6.
Let $S$ be a group, $H^{\prime}\subseteq S$ a subgroup of finite index and $N$ a $kS$-module. Assume that $\mbox{res}_{H^{\prime}}^{S}N\in{\tt Inj}(kH^{\prime})$ and $\mbox{res}_{S^{\prime}}^{S}N\in{\tt Inj}(kS^{\prime})$ for any finitely generated subgroup $S^{\prime}\subseteq S$. Then, $N\in{\tt Inj}(kS)$.
Proof. First of all, we shall prove that $\mbox{res}_{S^{\prime}}^{S}N\in{\tt Inj}(kS^{\prime})$ for any countable subgroup $S^{\prime}\subseteq S$. To that end, we fix a countable subgroup $S^{\prime}\subseteq S$ and express it as the union of an increasing sequence $(S_{i})_{i}$ of finitely generated subgroups. If $M$ is a $kS^{\prime}$-module, we consider the sequence of $kS^{\prime}$-modules $(M_{i})_{i}$, where $M_{i}=\mbox{ind}_{S_{i}}^{S^{\prime}}\mbox{res}_{S_{i}}^{S^{\prime}}M$ for all $i$. The inclusions $S_{i}\hookrightarrow S_{i+1}$ induce on $(M_{i})_{i}$ the structure of a direct system, whose colimit is $M$. Hence, there exists a short exact sequence of $kS^{\prime}$-modules
$0\longrightarrow{\textstyle{\bigoplus_{i}}}M_{i}\longrightarrow{\textstyle{\bigoplus_{i}}}M_{i}\longrightarrow M\longrightarrow 0$
Since $\mbox{Ext}^{n}_{kS^{\prime}}\!\left(M_{i},\mbox{res}_{S^{\prime}}^{S}N\right)\!=\mbox{Ext}^{n}_{kS_{i}}\!\left(\mbox{res}_{S_{i}}^{S^{\prime}}M,\mbox{res}_{S_{i}}^{S}N\right)=0$ for all $n>0$ and all $i$, the associated long exact sequence of Ext-groups that corresponds to the above short exact sequence of $kS^{\prime}$-modules implies that $\mbox{Ext}^{n}_{kS^{\prime}}\left(M,\mbox{res}_{S^{\prime}}^{S}N\right)=0$ for all $n>1$. This is the case for any $kS^{\prime}$-module $M$ and hence $\mbox{id}_{kS^{\prime}}\mbox{res}_{S^{\prime}}^{S}N\leq 1$. Then, Proposition 5.2(ii) implies that $\mbox{res}_{S^{\prime}}^{S}N\in{\tt Inj}(kS^{\prime})$.
We shall now prove that $N$ is an injective $kS$-module, using induction on the cardinality $\kappa$ of $S$. If $\kappa\leq\aleph_{0}$, then the result follows from the case considered above. If $\kappa$ is uncountable, then we may express $S$ as a continuous ascending union of subgroups $(S_{\lambda})_{\lambda<\kappa}$, each one having cardinality $<\kappa$. Our induction hypothesis implies that the $kS_{\lambda}$-module $\mbox{res}^{S}_{S_{\lambda}}N$ is injective for all $\lambda$. Let $M$ be a $kS$-module and define $M_{\lambda}=\mbox{ind}_{S_{\lambda}}^{S}\mbox{res}^{S}_{S_{\lambda}}M$ for all $\lambda$. We note that
$\mbox{Ext}^{n}_{kS}(M_{\lambda},N)=\mbox{Ext}^{n}_{kS_{\lambda}}\!\left(\mbox{res}^{S}_{S_{\lambda}}M,\mbox{res}_{S_{\lambda}}^{S}N\right)=0$
for all $\lambda$ and $n>0$. The ascending family of subgroups $(S_{\lambda})_{\lambda<\kappa}$ induces a continuous direct system of $kS$-modules $(M_{\lambda})_{\lambda<\kappa}$ with surjective structure maps, whose colimit is $M$. The short exact sequence of $kS$-modules
$0\longrightarrow K_{\lambda}\longrightarrow M_{0}\longrightarrow M_{\lambda}\longrightarrow 0,$
where $K_{\lambda}$ is the kernel of the structure map $M_{0}\longrightarrow M_{\lambda}$, is the $\lambda$-th term of a continuous direct system of short exact sequences. The colimit of the latter direct system of short exact sequences is the short exact sequence of $kS$-modules
(3) $0\longrightarrow K\longrightarrow M_{0}\longrightarrow M\longrightarrow 0$
Here, $K$ is equal to the continuous ascending union of its submodules $(K_{\lambda})_{\lambda<\kappa}$. Since $K_{\lambda+1}/K_{\lambda}$ can be identified with the kernel of the (surjective) structure map $M_{\lambda}\longrightarrow M_{\lambda+1}$, we conclude that $\mbox{Ext}^{n}_{kS}(K_{\lambda+1}/K_{\lambda},N)=0$ for all $\lambda$ and $n>0$. Then, Eklof’s lemma [24, Theorem 7.3.4] implies that $\mbox{Ext}^{n}_{kS}(K,N)=0$ for all $n>0$; hence, the short exact sequence (3) above implies that $\mbox{Ext}^{n}_{kS}(M,N)=0$ for all $n>1$. Since this is the case for any $kS$-module $M$, we conclude that $\mbox{id}_{kS}N\leq 1$. Using once more Proposition 5.2(ii), we conclude that $N$ is injective. $\Box$
Corollary 5.7.
(i) The class $\mathfrak{M}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$, $\Phi$ and $\Phi_{flat}$.
(ii) The class $\mathfrak{M}_{flat}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{flat}$.
(iii) The class $\mathfrak{M}_{inj}$ is closed under the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi_{inj}$.
Proof. Theorem 5.3 implies that any group contained in ${\scriptstyle{{\bf LH}}}\mathfrak{M}\cup\Phi\mathfrak{M}\cup\Phi_{flat}\mathfrak{M}$ satisfies Moore’s conjecture. In other words, we have ${\scriptstyle{{\bf LH}}}\mathfrak{M}\cup\Phi\mathfrak{M}\cup\Phi_{flat}\mathfrak{M}=\mathfrak{M}$, whence (i) holds. Analogously, Theorem 5.4 implies assertion (ii) and Theorem 5.5 implies assertion (iii). $\Box$
As a consequence of Chouinard’s work [14], we know that finite groups satisfy Moore’s conjecture. In fact, all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups satisfy Moore’s conjecture; see [1] for the original (projective) version of the conjecture and [2] for the injective version of the conjecture (which implies the flat version, in view of [2, Proposition 2.6]). Let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed. We can describe $\overline{\mathfrak{F}}$-groups hierarchically, as explained in §$\S 1$.IV. The following Corollary is precisely Theorem B of the Introduction; it is itself an immediate consequence of Corollary 5.7(i), in view of the inclusion $\mathfrak{F}\subseteq\mathfrak{M}$.
Corollary 5.8.
All $\overline{\mathfrak{F}}$-groups satisfy Moore’s conjecture over $k$, i.e. $\overline{\mathfrak{F}}\subseteq\mathfrak{M}$.
Remark 5.9. We may use Corollary 5.7(ii),(iii) and obtain analogues of Corollary 5.8 for the flat and the injective versions of Moore’s conjecture.
Acknowledgments. Wei Ren was supported by the National Natural Science Foundation of China (No. 11871125).